# Long-Range Pairing in the Kitaev Model: Krylov Subspace Signatures

### v2

@@ Abstract
Krylov subspace methods quantify operator growth in quantum many-body systems through Lanczos coefficients that encode how operators spread under time evolution. While these diagnostics have been proposed to distinguish quantum chaos from integrability, quadratic fermionic Hamiltonians are widely expected to exhibit trivial Lanczos structure. Here we demonstrate that Lanczos coefficients generated from local boundary operators provide a quantitative diagnostic of whether the lowest excitation gap is controlled by boundary-localized or bulk-extended modes in the long-range Kitaev chain, the model for topological superconductivity with algebraically decaying couplings. We introduce Krylov staggering parameter, defined as the logarithmic ratio of consecutive odd and even Lanczos coefficients, whose sign structure correlates robustly with the edge versus bulk character of the gap across the full phase diagram. This correlation arises from a bipartite Krylov structure induced by pairing, power-law couplings, and open boundaries. We derive an exact single-particle operator Lanczos algorithm that reduces the recursion from exponentially large operator space to a finite-dimensional linear problem, achieving machine precision for chains of hundreds of sites. These results establish Krylov diagnostics as operational probes of how low-energy excitations are localized along the chain and how strongly they are tied to the boundaries with broken U(1) symmetry, with potential applications to trapped-ion and cold-atom quantum simulators.
@@ I Introduction
Our main result is that the Lanczos coefficients provide an operational diagnostic of whether the lowest excitation scale is set by boundary-localized or bulk-extended modes in the long-range Kitaev chain. The key technical step is an exact single-particle Lanczos construction: for quadratic fermionic Hamiltonians the commutator algebra closes on operators linear in Majorana modes, so Heisenberg evolution of such operators is governed by a finite-dimensional Hermitian generator (single-particle Liouvillian) $L_{\mathrm{sp}}=i\mathcal{H}_{M}$. This reduces the Krylov problem from an exponentially large operator space to a $2N$-dimensional linear space, enabling machine-precision computation of Lanczos coefficients for chains with hundreds of sites.
On the physics side, we define an edge-dominated gap regime in which the smallest positive Bogoliubov-de Gennes (BdG) excitation is boundary localized, and a bulk-dominated gap regime in which it is extended in the bulk (Sec. IV). For Hermitian boundary seeds such as the edge Majorana $\gamma_{1}$, all diagonal Lanczos coefficients vanish ($a_{n}=0$), yielding a purely off-diagonal tridiagonal Krylov Hamiltonian (Appendix B.2). We then introduce and study the sequence of odd-even Krylov staggering parameters, defined as the logarithmic ratio of consecutive off-diagonal Lanczos coefficients $b_{n}$:
$\eta_{n}\equiv\ln\!\left(\frac{b_{2n-1}}{b_{2n}}\right),$ (1)
@@
and show that its sign structure correlates with edge- versus bulk-gap character across the long-range phase diagram. The underlying mechanism is rooted in a bipartite Krylov structure: for a real boundary seed and real antisymmetric Majorana Hamiltonian $\mathcal{H}_{M}$, successive Krylov vectors alternate between real and purely imaginary subspaces, so $L_{\mathrm{sp}}^{2}$ acts separately on odd and even sectors. When a low-energy edge mode dominates the seed overlap, the Krylov recursion produces an imbalance between the odd and even subsequences as recursion depth increases, leading to robust sign changes in $\eta_{n}$. When the lowest scale is set by bulk modes, the two subsequences remain more evenly matched and $\eta_{n}$ retains a fixed sign. While Ref. [13] studied topological transitions using state (spread) Krylov complexity, and Ref. [14] analyzed monitored Su–Schrieffer–Heeger (SSH) dynamics in Krylov space within a similar state-spread framework (including connections to quantum Fisher information), we instead use the operator Lanczos algorithm and derive an exact single-particle formulation to extract Lanczos coefficients for boundary seeds, which directly resolve the edge–bulk gap competition. In this way, operator dynamics captured in Krylov subspace provide a practical probe of boundary-versus-bulk gap control, despite the model being quadratic.
@@
The remainder of this paper is organized as follows. In Sec. II we introduce the long-range Kitaev Hamiltonian and its BdG and Majorana formulations, and we establish closure of the commutator algebra on Majorana-linear operators. Sec. III derives the single-particle operator Lanczos algorithm, derives the associated tridiagonal representation of the single-particle Liouvillian, and proves that the diagonal Lanczos coefficients vanish identically for Hermitian seed operators (Appendix B.2). In Sec. IV, we define edge- versus bulk-gap regimes and introduce the Krylov staggering parameter diagnostic. Sec. V presents Lanczos-coefficient data and joint phase diagrams showing quantitative agreement between the gap-based and Krylov-based classifications. Sec. VI discusses implications, including disordered long-range settings [17], interacting generalizations, and experimental prospects in cold-atom and trapped-ion platforms [50, 32].
@@ II.5 Liouvillian in the linear sector
III Lanczos Algorithm Derived for the Single-Particle Picture
@@
This section derives and develops the numerical construction of Lanczos algorithm and the associated Lanczos coefficients for the long-range Kitaev chain by exploiting the exact closure of Heisenberg dynamics on operators linear in Majorana modes. Accordingly, all Krylov dynamics reported here take place in the $2N$-dimensional coefficient space associated with operators linear in Majoranas.
@@ III.3 Lanczos recursion and tridiagonalization
The recursion terminates when $b_{n+1}$ falls below a numerical tolerance where we have implemented partial reorthogonalization. In practice we use $10^{-7}$, and the Krylov dimension is bounded by $\mathcal{K}\leq 2N$.
@@ III.4 Numerical implementation
For an open chain of $N$ fermionic sites, we construct the $2N\times 2N$ Majorana single-particle generator $\mathcal{H}_{M}$ appearing in Eq. (11), and hence the Hermitian single-particle Liouvillian matrix $L_{\mathrm{sp}}=i\mathcal{H}_{M}$ acting on the $2N$-component coefficient vector $u$ of operators linear in Majoranas. We then run the Lanczos recursion to obtain $\{b_{n}\}$ and the Krylov dimension $\mathcal{K}\leq 2N$, form the corresponding $\mathcal{K}\times\mathcal{K}$ tridiagonal matrix $T$, and evolve $\phi(t)$ by diagonalizing $T$. The recursion is terminated when the coefficient sequence becomes numerically unstable, and only Lanczos coefficients that satisfy multiple rigorous stability criteria are retained in the analysis; once any criterion is violated, that coefficient and all subsequent ones are excluded (see Appendix B.4 for details). Additional seed operators used to test robustness are listed in Appendix D.
@@ IV.1 Edge gap versus bulk gap
indicating that the lowest excitation is extended in the bulk. This classification provides an operational diagnostic for whether low energy physics is dominated by boundary or bulk degrees of freedom [60]. Numerical details and further discussions about $\omega_{\mathrm{edge}}$ and $\ell_{\mathrm{edge}}$ are provided in Section V.3 and data for alternative seed are available in Appendix D.

### v3

@@ Abstract
Krylov subspace methods quantify operator growth in quantum many-body systems through Lanczos coefficients that encode how operators spread under time evolution. Although these diagnostics were originally motivated by questions of chaos and integrability, quadratic fermionic Hamiltonians are often expected to exhibit trivial Lanczos structure. Here we show that, in the long-range Kitaev chain, Lanczos coefficients generated from local boundary operators sharply diagnose whether the lowest excitation gap is controlled by boundary-localized or bulk-extended modes. We introduce the Krylov staggering parameter for the Lanczos coefficients. In the short-range Kitaev chain with balanced hopping and pairing, we derive analytically for arbitrary system size (valid in the thermodynamic limit) and show that this quantity is exactly constant and its sign cleanly distinguishes the topological phase with Majorana edge modes from the trivial phase. Away from that limit, long-range couplings and pairing-hopping imbalance deform the simple flat structure and analytical control is lost, nevertheless, we show that the sign pattern of the diagnostic still tracks whether the lowest excitation gap is controlled by boundary modes or by bulk excitations. These results are enabled by an exact single-particle operator Lanczos algorithm, as derived in this work, which reduces the recursion from exponentially large operator space to a finite-dimensional linear problem and achieves machine precision for chains of hundreds of sites. Krylov diagnostics thus emerge as practical probes of boundary-versus-bulk low-energy physics in topological superconductors with broken $U(1)$ symmetry and algebraically decaying couplings.
@@ I Introduction
Our main results are twofold. First, for the short-range Kitaev chain at the hopping-pairing-balanced point, we derive an exact closed-form expression for the entire Lanczos recursion generated by the boundary Majorana seed $\gamma_{1}$: for every recursion depth $n$ and every system size $N$ (all the way to the thermodynamic limit), the off-diagonal coefficients $b_{n}$ are constant and can be obtained exactly, while all diagonal coefficients vanish, $a_{n}=0$ (Appendix C). Equivalently, the Krylov staggering parameter $\eta_{n}\equiv\ln(b_{2n-1}/b_{2n})$ is exactly $n$-independent, with no finite-size corrections and no approximations. This result is not merely a compact formula for Lanczos coefficients: the balanced short-range Kitaev chain in Majorana form is exactly the SSH chain, and the Krylov recursion generated by the boundary Majorana seed preserves this alternating-bond matrix structure. We find, the sign of $\eta_{n}$ is an exact analytic topological order parameter. In particular, $\eta_{n}<0$ if and only in the topological phase of the physical Kitaev chain with Majorana zero modes, while $\eta_{n}>0$ if and only in the trivial phase, thereby recovering Kitaev’s phase boundary from a purely Krylov-recursion argument without invoking band topology, winding numbers, or bulk-gap closure.
Second, for the long-range Kitaev chain, the exact flat short-range structure is deformed by power-law couplings, so $\eta_{n}$ is no longer constant in $n$. The crucial lesson from the short-range theorem, however, survives: each value of $\eta_{n}$ continues to encode the local SSH-type bonding character of the Krylov chain at recursion depth $n$, with $\eta_{n}<0$ corresponding to a locally topological strong-bond regime and $\eta_{n}>0$ to a locally trivial weak-bond regime. This motivates a new diagnostic principle for the deformed problem: not the constancy of $\eta_{n}$, but its sign structure across $n$. We show empirically across the full long-range/short-range and pairing imbalance phase diagram that sign changes of $\eta_{n}$ correlate robustly with regimes in which the lowest excitation gap is controlled by boundary-localized modes, whereas a fixed positive sign correlates with bulk-dominated trivial regimes. The key technical ingredient enabling both statements is an exact single-particle Lanczos construction: because the commutator algebra closes on operators linear in Majorana modes, Heisenberg evolution is governed by the finite-dimensional Hermitian generator $L_{\mathrm{sp}}=i\mathcal{H}_{M}$, reducing the Krylov problem from exponentially large operator space to a $2N$-dimensional linear problem and allowing machine-precision calculations for chains with hundreds of sites. In this way, Krylov subspace dynamics become a practical and quantitatively sharp probe of whether the low-energy gap is edge-dominated or bulk-dominated even in a quadratic, non-chaotic system.
@@
The underlying mechanism is rooted in a bipartite Krylov structure: for a real boundary seed and real antisymmetric Majorana Hamiltonian $\mathcal{H}_{M}$, successive Krylov vectors alternate between real and purely imaginary subspaces, so $L_{\mathrm{sp}}^{2}$ acts separately on odd and even sectors. When a low-energy edge mode dominates the seed overlap, the Krylov recursion produces an imbalance between the odd and even subsequences as recursion depth increases, which leads to robust sign changes in $\eta_{n}$. When the lowest scale is set by bulk modes, the two subsequences remain more evenly matched and $\eta_{n}$ retains a fixed sign. While Ref. [14] studied topological transitions using state (spread) Krylov complexity, and Ref. [15] analyzed monitored Su–Schrieffer–Heeger (SSH) dynamics in Krylov space within a similar state-spread framework (including connections to quantum Fisher information), we instead use the operator Lanczos algorithm and derive an exact single-particle formulation to extract Lanczos coefficients for boundary seeds, which directly resolve the edge–bulk gap competition. In this way, operator dynamics captured in Krylov subspace provide a practical probe of boundary-versus-bulk gap control, despite the model being quadratic.
@@
The remainder of this paper is organized as follows. In Sec. II we introduce the long-range Kitaev Hamiltonian and its BdG and Majorana formulations, establish closure of the commutator algebra on Majorana-linear operators, and derive the exact Krylov structure of the hopping-pairing-balanced short-range case in the thermodynamic limit (Sec. II.6), which serves as an analytic anchor for the long-range analysis. Sec. III derives the single-particle operator Lanczos algorithm, derives the associated tridiagonal representation of the single-particle Liouvillian, and proves that the diagonal Lanczos coefficients vanish identically for Hermitian seed operators (Appendix B.2). In Sec. IV, we define edge- versus bulk-gap regimes and introduce the Krylov staggering parameter diagnostic. Sec. V presents Lanczos-coefficient data and joint phase diagrams showing quantitative agreement between the gap-based and Krylov-based classifications. Sec. VI discusses implications, including disordered long-range settings [18], interacting generalizations, and experimental prospects in cold-atom and trapped-ion platforms [52, 33].
@@ II.5 Liouvillian in the linear sector
II.6 Exact Krylov structure of the short-range limit
The single-particle Liouvillian $L_{\mathrm{sp}}=i\mathcal{H}_{M}$ derived above admits a fully analytic treatment in the short-range limit $\alpha\to\infty$, $\epsilon=0$. In this limit only nearest-neighbor couplings survive,
Eq. (1) reduces to the standard Kitaev chain with hopping $t=\sin\theta$, pairing $\Delta=\sin\theta(1+\epsilon)$, and chemical potential $\mu=-2\cos\theta$ (explicit Hamiltonian in Appendix C, Eq. (61)). For analytically tractable $\epsilon=0$, we have the hopping-pairing balanced situation $\Delta=t$. The corresponding $H_{M}$ becomes a tridiagonal real antisymmetric matrix: acting on any Majorana mode $\gamma_{k}$, the Liouvillian $L_{\rm sp}$ couples it only to its two nearest neighbors k1k-1 and $k+1$, with alternating bond strengths $\mu$ intra-site and $2t$ inter-site. Equivalently, the balanced short-range Kitaev chain in its Majorana single-particle representation is exactly the SSH chain with intracell bond $\mu$ and intercell bond $2t$, and the Krylov recursion generated from $\gamma_{1}$ simply preserves this alternating-bond matrix structure step by step. As a direct consequence, the Krylov algorithm seeded with the boundary Majorana $\gamma_{1}$ traverses the Majorana chain sequentially, and the Lanczos off-diagonal coefficients simply read off those alternating bond strengths. The complete mathematical proof by strong induction is given in Appendix C.2; here we state the result.
Exact thermodynamic-limit result.
For the short-range Kitaev chain at $\Delta=t$ (i.e., $\epsilon=0$) where $\Delta$ and $t$ are the pairing and hopping terms respectively, with boundary seed $\gamma_{1}$, the Lanczos recursion produces, for every $n\geq 1$ and every system size $N\geq 1$:
$b_{2n-1}=|\mu|,\qquad b_{2n}=2|t|,\qquad a_{n}=0.$ (14)
Here $\mu$ is the chemical potential. See Eq. (61) in Appendix C where the relation with the notation used in Eq. (1) is established. There are no finite-size corrections, no transients, and no approximations. Equation (14) is exact for any $N$ and therefore remains exact in the thermodynamic limit as well.
Exact topological order parameter.
Define the Krylov staggering parameter as the logarithmic ratio of consecutive off-diagonal Lanczos coefficients:
$\boxed{\eta_{n}\equiv\ln\!\frac{b_{2n-1}}{b_{2n}}}.$ (15)
By Eq. (14), this quantity is exactly constant in both $n$ and $N$:
$\eta_{n}=\ln\!\frac{|\mu|}{2|t|}=\mathrm{const.},\quad\forall\,n,\,N.$ (16)
Its sign is an exact topological indicator:
$\mathrm{sgn}(\eta_{n})=\begin{cases}-1&|\mu|<2|t|\quad\text{(topological)},\\ \phantom{-}0&|\mu|=2|t|\quad\text{(critical point)},\\ +1&|\mu|>2|t|\quad\text{(trivial)},\end{cases}$ (17)
in exact agreement with Kitaev’s phase boundary [35, 2]. This is not a numerical observation or an asymptotic statement — it is a theorem: the sign of $\eta$ encodes the topological phase of the short-range Kitaev chain exactly, derived purely from the Krylov recursion algebra.
II.7 Krylov–SSH correspondence and topological order parameter
It is known in the literature [49, 44] that the Lanczos recursion maps operator dynamics onto a tight-binding problem on the Krylov chain, whose hopping amplitudes are the off-diagonal Lanczos coefficients $b_{n}$. In generic systems this chain is semi-infinite and structurally featureless; here we prove something sharply stronger and exactly: for the short-range Kitaev chain at $\Delta=t$, the balanced physical Kitaev chain in its Majorana single-particle representation is itself the paradigmatic SSH chain [58, 5], and the Krylov chain generated by the boundary Majorana seed is algebraically identical — as a matrix, not merely in spirit — to that same SSH chain. This identification is exact, holds for every system size $N$ (all the way up to the thermodynamic limit) and every recursion depth $n$, and is proved by strong induction in Appendix C.2C.3. It implies that the topological phase of the Kitaev chain and the topological phase of the Krylov chain are the same object, read off from the same mathematical structure: when $|\mu|<2|t|$, the Krylov chain is in the SSH topological phase, its left boundary site ($v_{0}\leftrightarrow\gamma_{1}$) is undimerized, and this is precisely the condition for Majorana zero modes in the physical chain.
The exact flat Lanczos structure of Eq. (14) admits a direct geometric interpretation: the Majorana matrix $\mathcal{H}_{M}$ at $\Delta=t$ is a real antisymmetric tridiagonal matrix whose entries alternate between $|\mu|$ (intra-site, $\gamma_{2j-1}$$\gamma_{2j}$) and $2|t|$ (inter-site, $\gamma_{2j}$$\gamma_{2j+1}$). Under the identification $A_{j}\leftrightarrow\gamma_{2j-1}$, $B_{j}\leftrightarrow\gamma_{2j}$, $t_{1}\leftrightarrow|\mu|$, $t_{2}\leftrightarrow 2|t|$, the single-particle Liouvillian $L_{\mathrm{sp}}=i\mathcal{H}_{M}$ is algebraically identical to the SSH Hamiltonian [58, 5] — not an analogy, but an exact matrix equality confirmed to machine precision (Appendix C.3). Because the induction proof of Appendix C.2 shows that the Krylov recursion traverses the Majorana chain one site at a time (Eq. (80)), the Lanczos off-diagonal coefficients $b_{n}$ are simply the bond strengths read off sequentially, and the Krylov tridiagonal is structurally the same alternating chain. The three objects — real-space Majorana chain, SSH Hamiltonian, and Krylov tridiagonal — are therefore in exact one-to-one correspondence (see the dictionary in Appendix C.3). The topological criterion follows immediately from SSH physics [58, 5]: the SSH chain is topological whenever the intercell bond dominates ($t_{2}>t_{1}$), which under the identification above is $2|t|>|\mu|$, i.e. $|\mu|<2|t|$ — exactly Kitaev’s phase boundary [35]. The Krylov staggering parameter $\eta_{n}=\ln(|\mu|/2|t|)$, being the logarithmic ratio of the weak (intracell) to strong (intercell) bond, is therefore an exact, $n$-independent, size-independent topological order parameter of the short-range Kitaev chain at $\epsilon=0$: its sign encodes the phase without any reference to bulk gap closure, winding numbers, or band topology (Appendix C.3, Eq. (102)). We emphasize the important caveat that this correspondence operates at the level of the single-particle Majorana matrix $L_{\mathrm{sp}}$, not the full many-body Hamiltonian; the Kitaev chain is a $p$-wave superconductor while the SSH chain is an insulator, and they differ in symmetry class and many-body structure. The precise statement — and its structural origin in the cancellation of all next-nearest-neighbor bonds at $\Delta=t$ — is detailed in Appendix C.3.
II.8 Motivation for the long-range study.
Equation (17) establishes $\eta_{n}$ as an analytically grounded, exactly proven topological diagnostic in the tractable short-range balanced limit: there, $\eta_{n}$ is constant in $n$ and its sign alone suffices to classify the phase. When long-range power-law couplings ($\alpha<\infty$) or a nonzero pairing imbalance ($\epsilon\neq 0$) are introduced, this flatness is broken — $\eta_{n}$ acquires a nontrivial dependence on recursion depth $n$ — and a constant-sign criterion no longer directly applies. The diagnostic must therefore be recast. The key insight, which the exact short-range result makes precise, is that each value of $\eta_{n}$ encodes the local Krylov SSH chain-type bonding character at recursion depth $n$: $\eta_{n}<0$ means the Krylov chain is locally in the topologically strong-bond regime at step $n$, and $\eta_{n}>0$ means it is locally in the trivially weak-bond regime. A sign flip of $\eta_{n}$ as a function of $n$ therefore signals that the Krylov recursion is traversing a crossover between these two regimes — the system’s low-energy structure is being probed at both topological and trivial energy scales as the recursion depth increases, a hallmark of a system in or near the topological phase. Conversely, when $\eta_{n}$ retains a fixed positive sign for all $n$, the Krylov chain remains uniformly in the trivial-bond regime throughout the recursion, signaling a bulk-dominated trivial phase. The remainder of this paper tests this hypothesis across the full $(\theta,\alpha)$ phase diagram of the long-range Kitaev chain at $\epsilon\neq 0$, and establishes empirically — motivated and anchored by the exact short-range proof — that the presence or absence of sign flips in $\eta_{n}$ as a function of $n$ provides a robust, operationally sharp diagnostic of the edge-versus-bulk character of the excitation gap, without requiring any reference to winding numbers, band topology, or bulk gap closure.
III Single-Particle Lanczos Algorithm
@@
This section derives and develops the numerical construction of the Lanczos algorithm and the associated Lanczos coefficients for the long-range Kitaev chain by exploiting the exact closure of Heisenberg dynamics on operators linear in Majorana modes. Accordingly, all Krylov dynamics reported here take place in the $2N$-dimensional coefficient space associated with operators linear in Majoranas.
The exact short-range result derived in Sec. II.6 provides a stringent benchmark for this construction. In the balanced short-range limit $\Delta=t$, with seed $\gamma_{1}$, the algorithm must reproduce $b_{2n-1}=|\mu|$, $b_{2n}=2|t|$, and $a_{n}=0$ for every recursion depth and every system size, up to machine precision (for a detailed proof, see Appendix C). The same construction is then used without modification in the long-range regime, where no closed-form expression is available and the Lanczos data become the central diagnostic.
@@ III.3 Lanczos recursion and tridiagonalization
The recursion terminates when $b_{n+1}$ falls below a numerical tolerance where we have implemented partial reorthogonalization. In practice we use $10^{-7}$, and the Krylov dimension is bounded by $\mathcal{K}\leq 2N$. Furthermore, to improve stability of the algorithm, we enforce symmetries that are shared by the Krylov vectors $v_{n}$ at every step, directly after $L_{sp}$ is applied to $v_{n}$. For example, a single Majorana initial seed, will always retain it’s parity order, meaning that all $v_{2n-1}$ will have the same parity as $v_{1}$ while $v_{2n}$ will have opposite parity.
@@ III.4 Numerical implementation
For an open chain of $N$ fermionic sites, we construct the $2N\times 2N$ Majorana single-particle generator $\mathcal{H}_{M}$ appearing in Eq. (10), and hence the Hermitian single-particle Liouvillian matrix $L_{\mathrm{sp}}=i\mathcal{H}_{M}$ acting on the $2N$-component coefficient vector $u$ of operators linear in Majoranas. We then run the Lanczos recursion to obtain $\{b_{n}\}$ and the Krylov dimension $\mathcal{K}\leq 2N$, form the corresponding $\mathcal{K}\times\mathcal{K}$ tridiagonal matrix $T$, and evolve $\phi(t)$ by diagonalizing $T$. The recursion is terminated when the coefficient sequence becomes numerically unstable, and only Lanczos coefficients that satisfy multiple rigorous stability criteria are retained in the analysis; once any criterion is violated, that coefficient and all subsequent ones are excluded (see Appendix B.4 for details). Additional seed operators used to test robustness are listed in Appendix E.
@@ IV Physics of Krylov Structure for Long-Range Pairing in Majorana Basis
The preceding section established an exact analytic anchor for everything that follows: in the hopping-pairing-balanced short-range limit, the Krylov staggering parameter $\eta_{n}\equiv\ln(b_{2n-1}/b_{2n})$ is an exact topological order parameter, constant in $n$ and fixed in sign by the phase of the Kitaev chain itself (Sec. II.6 and Appendix C). The purpose of the present section is to show how this exact short-range result can be operationally generalized to the long-range, pairing-imbalanced problem, where $\eta_{n}$ is no longer constant but its sign structure continues to encode whether the lowest excitation is boundary-dominated or bulk-dominated.
@@ IV.1 Edge gap versus bulk gap
indicating that the lowest excitation is extended in the bulk. This classification provides an operational diagnostic for whether low energy physics is dominated by boundary or bulk degrees of freedom [63]. Numerical details and further discussions about $\omega_{\mathrm{edge}}$ and $\ell_{\mathrm{edge}}$ are provided in Section V.3 and data for alternative seed are available in Appendix E.